OpenStudy (anonymous):

The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units?

OpenStudy (anonymous):

Perimeter is W + W + L + L, so enter in your values. Width is 11, length is 24 so; 24 + 24 + 11 + 11 = ?

OpenStudy (anonymous):

No, the rectangle cannot have x = 60 and y = 11 because x = 48 + 2y. No, the rectangle cannot have x = 60 and y = 11 because x = 24 + 2y. Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 48. Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 24. but then it gives me those to pick from

OpenStudy (anonymous):

So first, does the equation I showed you work? Remember 2w + 2l, or in this case 48 + 2y = 60 (y = 11)

OpenStudy (anonymous):

It doesn't right? So that means you eliminate C. and D.

OpenStudy (anonymous):

yeah it doesn't work

OpenStudy (anonymous):

So once you have that, you have to find your equation. So, 2w + 2l = P is the equation for perimeter. pretend you don't know the width, then fill in what you know.

OpenStudy (anonymous):

2w + 2(24) = 60 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now simplify that for me

OpenStudy (anonymous):

2w+48=60

OpenStudy (anonymous):

right, so now we just have to rearrange the terms a little bit, so we'll put 48 first, then use y for the width. 48 + 2y = 60

OpenStudy (anonymous):

Well, I guess your answer wants x = 48 + 2y. so what answer has that equation?

OpenStudy (anonymous):

y is 12?